OpenStudy (anonymous):

Solve x-2=sqrt x show steps

OpenStudy (anonymous):

\[x-2=\sqrt{x}\] \[(x-2)^2=x\] \[x^2-4x+4=x\] \[x^2-5x+4=0\] then just use quadratic equation.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

yw